# Multi-AUV Cooperative Underwater Multi-Target Tracking Based on Dynamic-Switching-enabled Multi-Agent Reinforcement Learning

### v2

@@ V-C Proposed cooperative tracking policy
Algorithm 3 FATHOM-Net-based DSBM tracking algorithm
@@ VI Evaluations
VI-A Simulation setup
The test is conducted on a computer equipped with an Intel(R) Core(TM) i9-12900H 2.50GHz processor and 32GB of RAM.
We construct a model using coordinate gridization, treating ET-AUVs as moving particles in a 3D underwater environment. The targets initially locate near the origin and move with a random velocity, while the ET-AUVs are uniformly distributed within a region one kilometer away from the targets.
The relevant evaluation notations is detailed in Table II.
TABLE II: Simulation notation
@@ VI-A Simulation setup
Figure 4: Convergence speed comparison under stable ocean currents (4 ET-AUVs tracking 2 targets)
Figure 5: Convergence speed comparison under ocean current interference (4 ET-AUVs tracking 2 targets)
Figure 6: Convergence speed comparison under stable ocean currents (12 ET-AUVs tracking 4 targets)
Figure 7: Convergence speed comparison under ocean current interference (12 ET-AUVs tracking 4 targets)
@@ VI-B Results
First, to measure the effectiveness of the proposed DSBM, we compare DSBM, with MADDPG [36], MAAC [37], and MADDPG-SAC [38]. Note that MADDPG-SAC is based on MADDPG where the entropy value (that is adaptively adjusted by SAC), is incorporated into the optimization goal to change the update of the Critic network. We compare DSBM with MADDPG-SAC to evaluate the significance of entropy value in the algorithm comparison. We conduct the comparison by respectively utilizing different algorithms to tracking the underwater targets in terms of convergence speed, tracking distance maintenance, difference in velocity from target, and intra-formation strategy consistency.
@@
Convergence speed: To evaluate the stability of algorithm convergence, we repeat multiple rounds of experiments (under different tracking scenarios) and select the samples in a 95% confidence interval. Totally, as shown in Fig. 4 to Fig. 7, our proposed DBSM performs the best than the other three compared objects in convergence speed.
Figure 8: Distribution of Agent-Target distance with stable ocean currents
Figure 9: Distribution of Agent-Target distance with ocean current interference
Figure 10: Agent-Target velocity difference with stable ocean currents
Figure 11: Agent-Target velocity difference with ocean current interference
In details, in Fig. 4, the proposed DSBM showcases the best convergence performance, even if the other algorithms eventually show a trend of convergence. It’s worth noting, MADDPG-SAC converges the fastest in the early stages of training, but the final convergence effect is not satisfactory. This is because the entropy value-based mechanism in MADDPG-SAC improves the exploration scope in the early stage of training, but it also results in that the training to be unstable in the later stages. Different with the case, as shown in Fig. 5, where the ocean current interference is employed, only the DSBM algorithm is able to converge. This demonstrates that our proposed DSBM’s superior adaptability to ocean currents.
Figure 12: Strategic consistency with stable ocean currents
Figure 13: Strategic consistency with ocean current interference
Furthermore, as shown in Fig. 6 and Fig. 7, DSBM also showcases the fastest convergence speed and the best performance with/without ocean current interference.
@@
The results in Fig. 4, Fig. 5, Fig. 6 and Fig. 7 demonstrate that: DSBM pays more attention to the information of better-performing ET-AUVs and conducts more meaningful exploration, thus the time the ineffective exploration can be significantly reduced.
@@
Tracking distance maintenance: As it is aforementioned, maintaining the distance between the agent and the target, i.e., the ”Agent-Target distance” during the tracking process is crucial. Thus, we test as well the distribution of Agent-Target distance with/without ocean current interference.
As the results shown in Fig. 8, with stable ocean current, DSBM can maintain the distance between ET-AUV and its targets within a small (pre-defined) distance change range. In contrast, the other algorithms all display significant deviations. Further, in Fig. 9, the similar results are also achieved in the scenario with ocean current interference.
@@
Velocity difference maintenance: In AUV swarm network-based underwater cooperative tracking, frequent acceleration and deceleration for ET-AUV can rapidly consume the energy, making against the long-duration tracking tasks especially in underwater. Therefore, the velocity difference between the ET-AUV and target, i.e., the ”Agent-Target velocity difference” can be used to measure the endurance capacity and tracking efficiency of the tracking policy.
As shown in Fig. 10, with stable ocean current conditions, the Agent-Target velocity differences of DSBM are within a very small range. In contrast, the other three algorithms showcase a wider range of Agent-Target velocity differences, indicating that the ET-AUVs have to frequently accelerate and decelerate throughout the entire tracking process. This will significantly reduce the endurance time under the same state conditions.
As shown in Fig. 11, with ocean current interference, the velocity of ET-AUVs guided by DSBM almost maintains consistency with the velocity of the tracking target. The results in both Fig. 10 and Fig. 11 demonstrate that the proposed DSBM can perform stable and high-efficient underwater cooperative tracking with complex underwater environment (with/without ocean current interference).
@@
Figure 14: Test for the availability of the proposed approaches in 3D environment.
Intra-formation strategy consistency: During the tracking process, a consistent strategy for the AUV formation helps improve efficiency in target identification and tracking, especially in complex underwater environments. This is crucial to prevent tracking errors or relief target loss due to environmental factors. Hence, we take the experimental scenario where twelve ET-AUVs tracking four targets as an example. And the ET-AUVs in each formation each choose actions from seven possible actions, to evaluate the consistency of these actions.
As the results shown in both Fig. 12 and 13, we showcase the probability distribution of the number of the consistent ET-AUVs in the formation. As shown in Fig. 12, before the ocean current interference is taken into account, the actions of all the ET-AUVs under our proposed DSBM are more likely consistent. Similarly, the proposed DSBM obviously performs the best than the other compared objects, while the ocean current interference is considered. This benefits from the proposed DSBM is based on our proposed FATHOM-Net framework, by which the AUV formation within a region undergoes local training, focusing more on the tracking targets within the region rather than the entire tracking task.
Finally, to make the entire tracking process directed by our proposed DSBM more clear, we use Unity to simulate the underwater environment for visualizing the tracking process. As shown in Fig. 14, the cubes represent the targets, and the spheres represent ET-AUVs. The red solid line represents the movement trajectory of the target, while other colored solid lines represent the tracking trajectories of the ET-AUVs. Specifically, we simulate three scenarios of multi-target tracking. Fig. 14(a)-Fig. 14(c) represent four ET-AUVs tracking two targets, Fig. 14(d)-Fig. 14(f) represent six ET-AUVs tracking three targets, and Fig. 14(g)-Fig. 14(i) represent twelve ET-AUVs tracking four targets. The visualization in Fig. 14 also proves the effectiveness of the proposed DSBM in tracking underwater targets in a 3D environment.
@@ VII Conclusion
In this study, we propose a dynamic-switching-enabled multi-agent reinforcement learning framework for performing AUV swarm-network-based cooperative multi-target tracking in underwater environment. We first give modelling approach for the underwater sonar-based detection and the ocean current interference. Then, we propose FATHOM-Net, a novel MARL architecture based on SDN, which enhances the flexibility and robustness of the AUV swarm. The proposed DSBM includes the ”dynamic-switching attention” and ”dynamic-switching resampling” mechanisms, both based on an experience replay buffer. It enhances AUV swarm network’s self-learning efficiency and accuracy. Finally, to achieve efficient formation classification in the interference of ocean currents, we propose an algorithm named ASMA. It efficiently and accurately assigns formations for ET-AUV through fuzzy logic and rule-based expert systems. Evaluation results demonstrate that our proposed tracking algorithm achieves the fastest convergence speed and tracking performance compared to various popular research products. Future research directions derived from this work can be summarized as the following:
  1. 1.
    Optimizing underwater obstacle avoidance mechanisms for ET-AUVs to mitigate potential damage;
  2. 2.
    Balancing energy consumption among the AUVs to enhance the endurance of the AUV swarm network-based cooperative tracking systems;
  3. 3.
    Designing AUV swarm robustness control framework when the unstable underwater communication (e.g., the underwater acoustic-based communication) has to been taken into account.

### v3

@@ V-C2 Proposed cooperative tracking algorithm
Algorithm 3 HSARL-based DSBM-driven tracking algorithm
@@ VI Evaluations
TABLE II: Simulation parameters
@@
$L_{e}$ Episode length 600
$B_{s}$ Buffer size 100000
$U_{i}$ Update interval 2000
$M_{s}$ Minimal buffer size to update 4000
$B$ Batch size 512
[…]
$\mu$ Fluid viscosity 1e-3 Pa·s
$D$ Damping factor 0.25
$\Delta t$ Simulation time step 0.1 s
VI-A Simulation setup
The test is conducted on a computer equipped with an Intel(R) Core(TM) i9-12900H, 2.50GHz processor and a RAM of 32GB.
We use MPE[36] as the simulation environment. Based on this,we construct a model using coordinate gridization, treating ET-AUVs as moving particles in a 3D underwater environment. The targets initially locate near the origin and move with a random velocity, while the ET-AUVs are uniformly distributed within a region one kilometer away from the targets.
We conduct simulations under three scenarios: 4 ET-AUVs tracking 2 targets, 6 ET-AUVs tracking 3 targets, and 12 ET-AUVs tracking 4 targets. Each scenario includes the conditions both with and without ocean current interference, respectively.
The relevant evaluation parameters are detailed in Table II.
(a) 4 ET-AUVs tracking 2 targets without ocean current interference
(b) 4 ET-AUVs tracking 2 targets with ocean current interference
(c) 12 ET-AUVs tracking 4 targets without ocean current interference
(d) 12 ET-AUVs tracking 4 targets with ocean current interference
(e) 4 ET-AUVs tracking 2 targets without ocean current interference
(f) 12 ET-AUVs tracking 4 targets without ocean current interference
Figure 4: Convergence speed comparison
@@ VI-B Results
First, to measure the effectiveness of the proposed DSBM, we compare DSBM, with MADDPG [37], MAAC [38], MADDPG-SAC [39], MATD3 [40], MAPPO [41], and MASAC [42]. Note that MADDPG-SAC is based on MADDPG where the entropy value (that is adaptively adjusted by SAC), is incorporated into the optimization goal to change the update of the Critic network. We compare DSBM with MADDPG-SAC to evaluate the significance of entropy value in the algorithm comparison. We conduct the comparison by respectively utilizing different algorithms to track the underwater targets in terms of convergence speed, tracking distance maintenance, velocity difference maintenance, intra-formation strategy consistency, energy consumption, tracing accuracy, respectively.
VI-B1 Convergence speed
@@
To evaluate the convergence of the proposed algorithm, we repeat multiple experiments in different tracking scenarios and select samples within a 95% confidence interval. Overall, as shown in Fig. 4a, Fig. 4b, Fig. 4c, and Fig. 4d, DBSM outperforms the other six methods in convergence speed.
Firstly, it should be noted, in Fig. 4a, Fig. 4b, and Fig. 4c, MATD3 clearly performs better than the other four algorithms, except for DBSM and MAPPO. This is due to MATD3’s noise injection mechanism. Adding noise to the policy network’s output (i.e., actions) encourages the agent to explore more in the state-action space. This helps the model avoid falling into local optima prematurely and gather more experience, leading to faster convergence.
Secondly, we note as well our algorithm occasionally falls into local optima but quickly recovered, as shown in the early stages of convergence line in Fig. 4d. This is due to that our proposed Dynamic-Switching Attention mechanism, which selects information from the best-performance ET-AUV in the current round, randomly chooses information from one or two other ET-AUVs, and then extracts features from the remaining ET-AUVs. This approach increases the proportion of information from the best-performance ET-AUVs in the training data, leading to better training results in similar scenarios and faster recovery from the early exploration phase. For more details about this, please refer to Proof A in Sec. V-C of this paper.
Thirdly, our proposed algorithm showcases the fastest training speed during the early stages, as displayed in Fig. 4a, Fig. 4b, Fig. 4c, and Fig. 4d. This is due to that our proposed Dynamic-Switching Resampling mechanism, which selects the top 50% of data based on reward values for updating. This help accelerate finding the optimal solution.
@@
In summary, the results in Fig. 4a, Fig. 4b, Fig. 4c, and Fig. 4d demonstrate that DBSM focuses more on the information from well-performing ET-AUVs, leading to more meaningful exploration and significantly reduces the time spent on ineffective exploration.
Additionally, to test the availability of the reward reshaping mechanism in DSBM, we conduct as well a convergence speed comparison between DSBM and DSBM without reward reshaping mechanism. After introducing the reward reshaping term, DSBM’s convergence improves significantly. As shown in Fig. 4e and Fig. 4f, the green curve represents the convergence under the proposed DSBM. It is obvious that the DSBM converges quickly than the DSBM (without reward reshaping mechanism) and is not easy to fall into a local optimum. That’s because, without reward reshaping mechanism, the tracking reward (in Eq.16) plays the most significant role in the reward function. The shorter the distance between the AUV and the target, the smaller the negative reward. However, in the early tracking stage, the distance between the AUV and the target is too large. As a result, AUV’s any action has a little effect on the system reward, making it easily falls into a local optimum state. After incorporating the reward reshaping mechanism, it encourages ET-AUV to move toward the target, especially in the early training stages. This effectively shortens the initial exploration phase, speeds up convergence, and prevents getting stuck in local optimum state.
(a) 12 ET-AUVs tracking 4 targets without ocean current interference
(b) 12 ET-AUVs tracking 4 targets with ocean current interference
Figure 5: Distribution of AUV-Target distance comparison
(a) 12 ET-AUVs tracking 4 targets without ocean current interference
(b) 12 ET-AUVs tracking 4 targets with ocean current interference
Figure 6: AUV-Target velocity difference comparison
VI-B2 Tracking distance maintenance
@@
As aforementioned, maintaining the distance between the ET-AUV and the target, i.e., the “AUV-target distance” is crucial during the tracking. Therefore, we also test the distribution of AUV-target distances with and without ocean current interference.
In conditions without ocean current interference, as shown in Fig. 5a, the ET-AUV guided by DSBM almost always maintains a distance of less than 0.5 from the target, outperforming the other models. This is because our model quickly achieves a stable tracking state in the early stages which significantly shortens the initial capture process and reducing outliers in the AUV-target distance.
When ocean current interference is introduced, as shown in Fig. 5b, the performance of the other six algorithms deteriorates to varying degrees, but the performance of DBSM improves. This is partly due to our reward reshaping term (Eq. 41), which gives higher rewards to AUVs moving toward the target, guiding them to maintain the correct direction.
Furthermore, the MAAC performs the second best, due to its layered attention mechanism. Specifically, the local attention layer in MAAC allows each AUV to focus on its own key information, while the global attention layer enables selective attention to important information from other AUVs. This means each AUV can effectively utilize other AUV’s information, ensuring each AUV maintains its distance from the target.
VI-B3 Velocity difference maintenance
@@
In AUV swarm network-based underwater cooperative target tracking, frequent acceleration and deceleration for each ET-AUV can rapidly consume the energy, making against the long-duration tracking tasks especially in underwater environment. Therefore, the velocity difference between the ET-AUV and target, i.e., the “AUV-Target velocity difference” can be used to measure the endurance capacity and tracking efficiency of the tracking policy.
As shown in Fig. 6a, the proposed DBSM performs the best with the smallest velocity difference without ocean current interference. MATD3 has the largest velocity difference, but over 80% of the differences are within 0.005. This indicates that although the velocity of AUVs under MATD3 vary greatly, the overall velocity difference is also small. This is because MATD3 introduces noise to the policy network’s output, which helps the model explore more in the state-action space and avoid local optima. However, it also increases the randomness in action selection, causing the AUV to occasionally deviate from the intended state, leading to minor velocity differences.
As shown in Fig. 6b, with ocean current interference, DBSM performs even better, with almost all data points close to the intended data (velocity). In contrast, the performance of the other algorithms declines to varying degrees. Although MATD3 shows a reduced maximum deviation, less than 40% of the points fall within a 0.005 difference, indicating an overall decline in performance.
TABLE III: Strategic consistency comparison (12 ET-AUVs tracking 4 targets without ocean current interference)
DSBM MADDPG-SAC MAAC MADDPG MATD3 MAPPO MASAC
All Different 41.15% 46.23% 46.12% 51.87% 43.12% 40.27% 44.63%
Two Alike 25.34% 28.06% 33.52% 34.22% 30.97% 31.45% 36.61%
All Alike 33.51% 25.71% 20.36% 13.91% 25.91% 28.28% 18.76%
TABLE IV: Strategic consistency comparison (12 ET-AUVs tracking 4 targets with ocean current interference)
DSBM MADDPG-SAC MAAC MADDPG MATD3 MAPPO MASAC
All Different 30.23% 48.79% 55.62% 53.10% 38.31% 50.12% 48.93%
Two Alike 18.25% 17.94% 17.38% 15.12% 24.11% 19.37% 26.98%
All Alike 51.52% 33.27% 27.00% 31.78% 37.58% 30.51% 24.09%
VI-B4 Intra-formation strategy consistency
During the tracking process, a consistent strategy for the ET-AUV formation helps improve target identification and tracking efficiency, especially in complex underwater environments. This is crucial to prevent tracking errors or relief target loss due to environmental factors.
We analyze the strategy consistency in a scenario where 12 ET-AUVs track 4 targets, with 3 ET-AUVs in each formation. There are three possible cases for strategy selections within each formation:
  1. 1.
    All Different: Each ET-AUV in the formation has a different strategy.
  2. 2.
    Two Alike: In the formation, there are two ET-AUVs whose strategies are the same.
  3. 3.
    All Alike: The strategy is the same for all ET-AUVs in the formation.
As shown in Table III, the proportion of identical strategies within the DSBM-guided formation is the highest, indicating the best tracking performance of DSBM. This is because the formation under DSBM reaches a stable tracking state quickly. At this stage, the ET-AUVs and the target form a shape similar to a triangular pyramid, where the three ET-AUVs form the base triangle, and the target serves as the apex. In this situation, the target’s position appears nearly identical from each AUV’s perspective, leading to a convergence in strategy selection within the formation. Therefore, DSBM-guided formation proportions of exhibit the highest strategy consistency.
As shown in Table IV, when ocean current interference is applied, the proportion of strategy consistency increases for all the algorithms. This is because, in our design, the interference from ocean currents is similar within close position. This means that each ET-AUV within the same formation resists the current’s force, resulting in convergent action strategy selection. Additionally, our algorithm shows the greatest improvement and ultimately achieves the highest strategy consistency. The above phenomenon results from the following factors: 1) DSBM-guided formations can achieve a stable state more quickly; 2) our proposed Dynamic-Switching mechanism strengthens the ET-AUVs’ resistance to ocean currents. Totally, both of the above factors result in consistent strategy selection when the current’s magnitude and direction are similar.
VI-B5 Energy consumption
To visually demonstrate DSBM’s contribution to improving the energy efficiency for the tracking system, we test the system remaining energy during the training. The system energy consumption follows the following rules: 1) for target tracking, the ET-AUV has seven limited actions, i.e., up, down, left, right, forward, backward, and staying still; 2) energy is consumed only when the ET-AUV moves; 3) staying still consumes no energy. As shown in Fig. 7a and Fig. 7b, DSBM achieves the most stable and efficient state. This is because our proposed DSBM encourages ET-AUVs to focus on velocity consistency. When the ocean current force aligns with the ET-AUV’s movement direction, the ET-AUV is more likely to choose the staying still action.
(a) 12 ET-AUVs tracking 4 targets without ocean current interference
(b) 12 ET-AUVs tracking 4 targets with ocean current interference
Figure 7: Energy consumption comparison
VI-B6 Tracking accuracy
In underwater cooperative tracking operations, tracking accuracy is a key factor in evaluating the model performance. We assume that if the actual distance between the ET-AUV and the target is within 0.005 of the preset distance, it counts as accurate tracking.
As shown in Table V, with ocean current interference, DSBM achieves a tracking accuracy of 94.23%. When the ocean current interference is introduced, DSBM’s tracking accuracy reaches 99.90%. In both cases, DSBM performs the best.
TABLE V: Tracking accuracy comparison (12 ET-AUVs tracking 4 targets)
DSBM MADDPG-SAC MAAC MADDPG MATD3 MAPPO MASAC
Without ocean current interference 94.23% 49.83% 55.40% 50.30% 83.30% 88.40% 46.83%
With ocean current interference 99.90% 28.77% 27.67% 26.77% 38.37% 41.97% 38.47%
@@ VI-B6 Tracking accuracy
Figure 8: Test for the availability of the proposed approaches in 3D environment (the word trajec. in the legend of each sub-figure is short for trajectory).
Finally, to make the entire tracking process directed by our proposed DSBM more clear, we use Unity [43] to simulate the underwater environment for visualizing the tracking process. We have embedded ocean currents and other marine elements into the MPE environment. It can realistically simulate our proposed method for multi-AUV tracking in underwater scenarios. Finally, we use Unity to visualize the trajectory data of each ET-AUV and the target. As shown in Fig. 8, the red cubes represent the targets, and the blue spheres represent ET-AUVs. The red solid line represents the movement trajectory of the target, while other colored solid lines represent the tracking trajectories of the ET-AUVs. Specifically, we simulate three scenarios of multi-target tracking. Fig. 8a-Fig. 8c represent 4 ET-AUVs tracking 2 targets, Fig. 8d-Fig. 8f represent 6 ET-AUVs tracking 3 targets, and Fig. 8g-Fig. 8i represent 12 ET-AUVs tracking 4 targets. The visualization in Fig. 8 also proves the effectiveness of the proposed DSBM in tracking underwater targets in a 3D environment.
@@ VII Conclusion
In this paper, we take a deep study on multi-AUV cooperative underwater multi-target tracking. We regard the AUV swarm as an underwater ad-hoc network and employ SDN technique to build HSARL architecture, a software-defined multi-AUV reinforcement learning architecture. Based on HSARL, we propose the Dynamic-Switching-based MARL (DSBM)-driven tracking algorithm. It includes ”Dynamic-Switching Attention” and ”Dynamic-Switching Resampling” mechanisms to improve the AUV swarm network’s self-learning efficiency and enhance tracking accuracy. Additionally, to further expedite the convergence of the HSARL algorithm, especially during the early learning phase, we introduce a reward reshaping mechanism to ensure more rapid and stable learning outcomes. Finally, we propose ASMA, an advanced classification algorithm to efficiently assigns formations in the presence of ocean current interference. Evaluation results demonstrate that our proposed tracking algorithm achieves the fastest convergence speed and tracking performance compared to various popular research products. Future research directions derived from this work can be summarized as the following: 1) Optimizing underwater obstacle avoidance mechanisms for ET-AUVs to mitigate potential damage; 2) Balancing energy consumption among the AUVs to enhance the endurance of the AUV swarm network-based cooperative tracking systems; 3) Designing AUV swarm robustness control framework when the unstable underwater communication (e.g., the underwater acoustic-based communication) has to been taken into account.
VIII ACKNOWLEDGEMENTS
The work is supported by the Joint Funds of the National Natural Science Foundation of China (No. U22A2011), Frontier Technologies R&D Program of Jiangsu (No. BF2024072), and National Training Program of Innovation and Entrepreneurship for Undergraduates (No. 202410145004).